# Singapo: Single Image Controlled Generation of Articulated Parts in Objects

### v2

@@ 5 Conclusion and Future Work
We address the problem of reconstructing household articulated objects from a single image in a generative manner. We break down this complex task into sub-tasks solving the problem from coarse to fine: we first infer part connectivity into a graph from the image, and then use the extracted graph and input image as guidance to produce part arrangement and motion parameters to describe the parts abstractly, and we finally retrieve the mesh for each part to assemble the full object. We develop a diffusion model with attention mechanism to distill DINOv2 features of the image and to leverage graph together to condition the part generation. We demonstrate the effectiveness and generalizabilty of our method on PartNet-Mobility and ACD datasets, and show that our method outperforms the state-of-the-art methods.
@@ Acknowledgments
This work was funded in part by CIFAR AI Chair, a Canada Research Chair, NSERC Discovery, and enabled by support from WestGrid and the Digital Research Alliance of Canada. We thank Han-hung Lee, Hou In Ivan Tam, Qirui Wu, and Yiming Zhang for helpful discussions.
@@ A.5 More Experiment Results
More qualitative results. We provide more qualitative results showing generated objects in animation for our model and other baselines in the supplementary video. The video shows 6 examples each for the PartNet-Mobility, ACD dataset, and the real images from the internet. We observe that our model can generalize well to the unseen and real images of objects with variational appearances in terms of part arrangement and object texture.

### v3

@@ 5 Conclusion and Future Work
We address the problem of reconstructing household articulated objects from a single image in a generative manner. We break down this complex task into sub-tasks solving the problem from coarse to fine: we first infer part connectivity into a graph from the image, and then use the extracted graph and input image as guidance to produce part arrangement and motion parameters to describe the parts abstractly, and we finally retrieve the mesh for each part to assemble the full object. We develop a diffusion model with attention mechanisms to distill DINOv2 features of the image and to leverage graphs together to condition the part generation. We demonstrate the effectiveness and generalizability of our method on PartNet-Mobility and ACD datasets, and show that our method outperforms the state-of-the-art methods.
@@ Acknowledgments
This work was funded in part by a CIFAR AI Chair, a Canada Research Chair, and NSERC Discovery grants, and enabled by support from the Digital Research Alliance of Canada. We thank Han-Hung Lee, Hou In Ivan Tam, Qirui Wu, and Yiming Zhang for helpful discussions.
@@ A.5 More Experiment Results
More qualitative results. We show more examples of our method tested on real images from the internet in Figure 10. We observe that our method can robustly generate plausible objects from real images with variations in appearance in terms of part arrangement and object texture, and even in the presence of backgrounds (as shown in the last two rows). We provide more qualitative results showing generated objects in animation for our model and other baselines in the supplementary video. The video shows 6 examples each for the PartNet-Mobility, ACD dataset, and the real images. We observe that our model can generalize well to the unseen and real images of objects with variational appearances in terms of part arrangement and object texture.
Figure 10: Examples of our method tested on real images from the internet. Our method can robustly generate plausible objects from real images with varied appearance in terms of the part arrangement and object texture, and also when there is a background in the image (see last two rows).
Editability of the generated objects. We design a staged approach to enable human intervention and to incorporate user changes for interactive generation. Figure 11 demonstrates the editability of the generated results by showing the process of modifying the graph and specifying the part attributes in the graph. We observe that our model can generate plausible objects that are consistent with the user’s intention. Even when the user-specified graph asks the model to generate fewer or more parts than the image suggests, our model can leverage the learned prior of articulated objects to make the necessary adjustments to the object to respect the user’s intention while maintaining object plausibility.
Test on images with objects in arbitrary state. Our model is trained on images of objects in a resting state. To evaluate the generalization of our model on unseen states, we test our model on images of objects in arbitrary states in a zero-shot setting. Specifically, we feed the model with renderings of the testing objects in the PartNet-Mobility dataset by randomizing articulated states and viewpoints. We show the qualitative results in Figure 12. We observe that our model is still able to generate plausible objects when the parts are with reasonable amount of articulation and visibility, as shown in the cases in the middle column. However, our method struggles to produce consistent results with the input image when the parts are in extreme states (e.g., with doors widely opened), as shown in the cases in the right column. The graph prediction from GPT-4o can be inaccurate in these cases due to the existence of self-occlusion. Our generative module presents difficulties in understanding the spatial structure of the parts under large differences in object shape from the training data. The revealed interior of the objects (e.g., the shelves inside the cabinet) can also be a distraction for the model. We also observe that our model cannot produce the objects with the same articulation as the image shows, which can be an interesting direction for future work.
Figure 11: Examples demonstrating how our method enables interactive generation by user-driven graph correction (see rows 1 and 2) or graph editing (see rows 3 and 4) and specifying the part attributes in the graph (see rows 5 and 6). Our model generates plausible objects that are consistent with the user’s intention. Even when the user-specified graph asks the model to generate fewer or more parts than the image suggests (as shown in rows 3 and 4), our model can make the necessary adjustments to the object to respect the user’s intention while maintaining plausibility.
Figure 12: Examples of our method tested on images of objects with arbitrary articulated states from random views. We observe that our model is still able to generate plausible objects when the parts are with reasonable amount of articulation and visibility, as shown in the cases in the middle column. However, our method struggles to produce consistent results with the input image when the parts are in extreme states (e.g., with doors widely opened), as shown in the cases in the right column.